OpenStudy (anonymous):

Find the sum of all 3-digit positive numbers N that satisfy the condition that the digit sum of N is 3 times the digit sum of N+3.

OpenStudy (anonymous):

i found 117,120 ; 207,210

OpenStudy (anonymous):

to be 2 of them ant more?

OpenStudy (anonymous):

Very interesting question.

OpenStudy (anonymous):

Suppose you have \[ N =d_1d_2d_3\\ N+3 =d_1d_2d_4 \\ \]Seems like in this case it won't work..

OpenStudy (shubhamsrg):

We'll have to take cases maybe . Let the number be a b c 1st case : c<=6 then, a+b+c = 3(a+b+c+3) 2(a+b+c) = -9 so no solution case 2: c=7 (a+b+7) = 3( a + (b+1) + 0 ) 4 = 2(a+b) a+b = 2 (a,b,c) can be (1,1,7) and (2,0,7) case 3: c=8 (a+b+8) = 3(a+ (b+1) + 1) 2 = 2(a+b) a+b=1 (a,b,c) can be (1,0,8) only. case 4: c=9 (a+b+9) = 3(a+ (b+1) +2 ) 3 = 2(a+b) No solutions 120 and 210 are not solutions! how did you write em? o.O

OpenStudy (shubhamsrg):

Only solutions thus are 108,117 and 207 in case I missed something ?

OpenStudy (anonymous):

108, 117 and 207 are the only solutions, you didn't miss anything

OpenStudy (anonymous):

Nice!!

OpenStudy (anonymous):

why did u take \[c \le 6\]

OpenStudy (anonymous):

and how did u get a+b+c = 3(a+b+c+3)??

OpenStudy (anonymous):

@shubhamsrg ??

OpenStudy (shubhamsrg):

for c>6, the last digit does not remain <9 , hence I took a case for c<6. and (a+b+c) = 3(a+b+ (c+3)) , that is only what the question says.

OpenStudy (anonymous):

\[\huge c=7=>a+b+c=3(a+b+10)\]

OpenStudy (anonymous):

\[\huge 2(a+b)=-23\]

OpenStudy (anonymous):

@shubhamsrg ??

OpenStudy (shubhamsrg):

at c= 7, on addition of 3, last digit becomes 0 and b gets succeeded by 1. hence new eqn will be a+b+ 7 = 3 ( a+ (b+1) + 0 )

OpenStudy (anonymous):

\[\huge thanks\]